# Geometry-Aware Diffusion Models for Multiview Scene Inpainting

### v1

@@ 3.2.2 Uncertainty-aware Multireference Conditioning
Using our cue-based notation, our method inpaints a single image via $\widehat{I}=f(I,M|A_{\mathcal{R}},\mathcal{G}_{\mathcal{R}})$, based on a reference set $\mathcal{R}$. However, fusing information across references is unclear, especially when conflicting information is present. We implement this by running parallel per-reference diffusion processes and fusing the output noise estimates at each diffusion step. Ideally, this fusion is geometry-aware (e.g., front-faces of $S_{i}$ provide high-certainty photometric information, while back-faces only upper bound target view depth). Since geometric maps cannot be directly used at inference due to errors in estimated geometry, we alter the diffusion model to estimate confidence maps for each reference, emulating the necessary geometric masks. In the following, we describe how our inpainting is actually implemented, including confidence estimation and fusion across parallel diffusion processes.
@@
Hierarchical Confidence Estimation. The geometric cues, $\mathcal{G}_{\mathcal{R}}$, indicate which areas of each reference are trustworthy. We utilize three geometric signals. First, the front-facing confidence mask, $\mathbf{C}_{f}$, indicates a pixel is either outside the inpainting mask in the target view or guided by a front-facing rendered pixel. In other words, the model has copied the photometric content from the target image itself or the rendered photometric content ($T_{i}(I_{i})$). Second, the back-facing confidence mask, $\mathbf{C}_{b}$, indicates a pixel is geometrically guided by a rendered back-face, suggesting new geometry must be generated in front of it. Finally, the shadow confidence mask, $\mathbf{C}_{s}$, signals the model’s certainty in trusting photometric information despite potential occlusion (shadow volume). Notice the confidence masks are closely related to the geometric cues $F_{i}$, $B_{i}$, and $C_{i}$. Importantly, though, the rendered maps are often slightly misaligned with the actual target image due to geometry estimation errors. Thus, we instead modify our diffusion model to estimate these three confidence masks, i.e., at every diffusion step, our inpainter not only generates a denoising estimate, but also provides an estimate of these three confidence masks. Please see §B.6 for details on supervising the confidence masks, and Fig. 7 for a visualization.
@@
Parallel Diffusion Processing. We now formalize the inpainting process. Given $\mathcal{R}$ and our diffusion model, $f$, we split the inpainter into $n=|\mathcal{R}|$ independent streams. Let $(\varepsilon_{i,t},\mathbf{C}_{f,i,t},\mathbf{C}_{b,i,t},\mathbf{C}_{s,i,t})=f_{i}(I,M|A_{i},\mathcal{G}_{i})$, where $\varepsilon_{i,t}$ is the estimated noise for reference $i$, $A_{i}=\{H,T_{i}(I_{i})\}$, and $\mathcal{G}_{i}=\{F_{i},B_{i},\widehat{D}_{i},C_{i}\}$, be the output of the $i$th process at diffusion time $t$. Denote $\mathcal{C}_{t}=\{\mathbf{C}_{f,i,t},\mathbf{C}_{b,i,t},\mathbf{C}_{s,i,t}\}_{i}$ as the combined confidence information for step $t$. We then fuse the noise estimates at time $t$, $\mathcal{E}_{t}=\{\varepsilon_{1},\ldots,\varepsilon_{n}\}$, to obtain a fused estimate, $\varepsilon_{t}=\Gamma(\mathcal{E}_{t},\mathcal{C}_{t})$. Before the next timestep, $t-1$, $\varepsilon_{t}$ is used to update all the diffusion processes across reference streams. This fuses multiview information in the learned, generative space of the diffusion noise estimate, rather than in pixel space. The fusion operator, $\Gamma$, follows a simple rules-based hierarchy, which we describe in §3.3.1.
@@ 3.2.3 Multiview-aware Training
We initialize our reference-guided inpainter as a Stable Diffusion v2, fine-tuned for inpainting [53, 60]. To condition on the cues, as in prior work [4], we add zero-initialized channels to the first convolutional layer of the diffusion UNet. We also modify the output of the noise estimator to output the confidence masks, alongside the noise estimate at each time step. For training, one can utilize a multiview dataset (real or synthetic); however, to ensure greater data diversity, we also synthesize data from single-view images, via monocular depth estimation, perturbing the virtual camera, and rendering the starting image as a reference. To simulate geometry estimation errors, we artificially perturb the reference mesh; in that case, supervision for the confidence masks can be generated via the unperturbed mesh. Additional details are provided in §B.6.
@@ 3.3 Autoregressive Scene Inpainting
With our geometry estimator (§3.1) and reference-aware view inpainter (§3.2), we can now iteratively inpaint the entire scene. To begin, we initialize the scene geometry by computing multi-view metric depth maps of the incomplete input views. We also initialize the “autoregressive set” with the incomplete input images, which will be autoregressively inpainted. A random view is then selected to start the inpainting. Each iteration consists of three steps: (i) geometry-aware inpainting of a subset of images that have not yet been inpainted, (ii) updating the autoregressive set and the scene geometry with the inpainted images, and (iii) selecting a subset of images that remain to be inpainted for the next autoregressive iteration. A high-level illustration of these steps is provided in Fig. 2 (b).
@@ 3.3.1 Geometry-aware Inpainting
For each target view, we select a subset of reference views from the autoregressive set, preferably those that have already been inpainted. Following §3.2, we render the appearance and geometric cues using all reference views, then run parallel diffusion processes across the references, fusing the noise prediction at each step via the estimated confidences. The fusion, $\Gamma$, involves four hierarchical confidence levels: (i) front-face confidence, (ii) backface confidence, (iii) shadow confidence, and (iv) no confidence. For each patch at each level, we pick the noise estimate from the closest camera among the views at the same level. This approach allows us to fuse the most reliable information from different reference images during the denoising diffusion process. Please see §B.5 for details.
@@ 3.3.2 Autoregressive Set and Scene Geometry Update
In this step, we replace each target view in the autoregressive set with its inpainted version, and update the geometry of the target views using the corresponding inpainted images, as elaborated in §B.1.
@@ 3.3.3 Selecting the Next Images to Inpaint
We employ a two-stage strategy for the autoregressive process. First, we inpaint a wide-baseline subset of the scene, one by one, to generate the missing content of the scene. Then, we propagate the generated content to the remaining views simultaneously. At the start of the first stage, we use a greedy min-max approach to select the wide-baseline subset. Beginning with the first view, we iteratively choose the view that maximizes the minimum distance to the views already in the subset. After selecting the wide-baseline subset, we arrange them so that each view has the minimum distance to the previous views (see §B.7 for details). The wide-baseline stage is performed in this sorted order. In this stage, each target view is conditioned on all views of the autoregressive set, either inpainted or not. In the propagation step, the target views (i.e., all remaining views) are only conditioned on the inpainted wide-baseline images.
@@ 4.1 Implementation Details
We use PyTorch3D [50] for mesh rendering. For training, we use a mixture of two datasets, MS COCO [31] and Google Scanned Objects [15]. For MS COCO, we synthesize the reference-based geometric cues using DepthAnything V2 [76]. Training and evaluation is performed on NVIDIA L40 GPUs (16 for training). Additional implementation details are provided in §C.
@@ 4.2 Evaluation Protocol
Datasets. To evaluate our method for narrow-baseline inpainting, we use the SPIn-NeRF [42] dataset, a widely used benchmark for object removal in front-facing real-world scenes. This dataset consists of 10 scenes, each containing 60 images that include the object to be removed and corresponding object masks. Each scene also includes 40 images without the object as ground truth for evaluation. In addition, we also use the scene-centric portion of the NeRFiller [72] dataset to further investigate the performance of our method on more complex scenes, particularly those with larger baselines. This dataset serves as a benchmark for the scene completion task. For the few-view inpainting task, we use SPIn-NeRF and the scene-centric portion of NeRFiller, resulting in a total of 15 scenes. For each scene, we uniformly sample eight subsets of two views and eight subsets of three views. This yields a total of 240 few-view sets.
@@
Baselines. For object removal, we compare our method to state-of-the-art approaches on the SPIn-NeRF dataset, specifically: SPIn-NeRF [42], which inpaints images and depth maps to supervise NeRF fitting; Inpaint3D [48], which uses SDS [47] to inpaint a NeRF with a diffusion prior; RefFusion [44], which adapts an inpainting diffusion model to a reference image and uses SDS to inpaint a 3DGS [25] with the reference-adapted diffusion model; InFusion [34], which inpaints a 3DGS by inpainting the depth map of an inpainted reference image; MVIP-NeRF [8], which uses SDS to inpaint a NeRF by inpainting the rendered images and normal maps via a diffusion prior, and MALD-NeRF [30], which inpaints a NeRF by performing masked adversarial training for per-scene customization of a diffusion model. InFusion [34] also evaluates Gaussian Grouping [77] on SPIn-NeRF, a 3DGS-based segmentation method that does not directly evaluate on SPIn-NeRF itself. For scene completion, we compare our method to NeRFiller [72], which alternates between editing input images, and updating the 3D representation encompassed by a NeRF to fuse the edited images, this process is collectively called Dataset Update. Finally, for the few-view task, we compare our method to SPIn-NeRF [42] and NeRFiller [72]. For this task, we use NeRFiller directly. However, since SPIn-NeRF relies on supervision via COLMAP’s [55] sparse depth, we disable this supervision for the few-view task, as scenes from the NeRFiller dataset lack COLMAP information.
@@
Figure 3: Qualitative comparisons on the SPIn-NeRF dataset. Notice other methods produce blurry regions (e.g., bench end) due to multiview inconsistencies, while ours preserves sharpness and visual plausibility. Please zoom in for details.
@@
Metrics. For the SPIn-NeRF dataset, we use the same evaluation protocol as reported in the original paper [42]. Specifically, we compute LPIPS [82] (with VGG-16 [58]) and FID [19] between the inpainted images and the ground-truth images, cropped by the inpainting mask’s bounding box. The bounding box’s size is increased by 10% before cropping, uniformly in each direction. We additionally assess the sharpness of the inpainted images within the inpainting mask using the Laplacian variance [46]. As our inpainting method does not explicitly enforce 3D consistency via a 3D radiance field, we evaluate our consistency using the TSED metric [78]. Here, our feature correspondences are limited to those inside the inpainting masks. On the SPIn-NeRF dataset, as the scenes have very small baselines, correspondences are considered across all possible view pairings.
@@
Finally, for the few-view inpainting task, as there is no ground truth available, we only compute sharpness and MUSIQ to evaluate image quality. We also compute Corrs and TSED on all possible image pairs in the few-view set. Unlike NeRFiller [72], we compute MUSIQ and Corrs only inside the bounding box around the inpainting mask for the few-view inpainting task. §E provides additional details on computing TSED.
@@ 4.3 Results
Object Removal on Narrow-baseline Scenes. Due to the narrow baseline, we perform single-reference inpainting for this task. As our method does not rely on fitting a NeRF on the training views, and the evaluations are performed on the test views, we follow SPIn-NeRF’s [42] procedure to evaluate image inpainters, e.g., LaMa [61]. Specifically, we first fit a NeRF on the training views and render the test views, which will now contain the unwanted object. The rendered test views are then used as inputs to our inpainting method. To evaluate SPIn-NeRF, InFusion, and MVIP-NeRF, we run their official code to reproduce the results. Since RefFusion and Inpaint3D do not provide publicly available code, we only show the numbers provided by the papers. For MALD-NeRF, we evaluate their publicly available inpainted images for the SPIn-NeRF dataset. As shown in Tab. 1, our method outperforms all baselines on the SPIn-NeRF benchmark. We obtain comparable sharpness and LPIPS to MALD-NeRF, but significantly outperform it on other metrics. We also demonstrate that our method achieves faster scene inpainting compared to other methods. Furthermore, the TSED results show that despite other methods utilizing an explicit or implicit radiance field to enforce 3D consistency, our method achieves a higher consistency score, primarily due to blurry outputs. Please see §G for a comprehensive TSED analysis. Finally, we present a set of qualitative results in Fig. 3, showing our method produces sharper images than baselines.
@@
Table 1: Quantitative evaluation of the object removal task on SPIn-NeRF dataset. We denote sharpness (×$\times 10^{-5}$) as $\sigma$, the percentage of consistent image pairs (TSED) at $T_{\text{error}}=2.0$px as $T_{2\text{px}}$, and average run time per scene as $\tau$. The first and second sections report numbers from [48] and [34], respectively, possibly using different evaluation code than SPIn-NeRF. The last section ensures direct comparability with consistent evaluation code; see §G for a comprehensive TSED analysis.
Figure 4: Qualitative comparisons on the NeRFiller dataset. NeRFiller can converge to blurry content, due to mixing divergent views, while ours generates and then propagates sharp content (e.g., see details in the backpack or window glass in the zoomed patches).
@@
Table 2: Evaluation of scene completion on the NeRFiller dataset. We denote sharpness (×$\times 10^{-5}$) as $\sigma$, the percentage of consistent image pairs (TSED) at $T_{\text{error}}=2.0$px as $T_{2\text{px}}$, and average run time per scene as $\tau$. $\cdot^{D}$: direct outputs of the inpainting model - $\cdot^{N}$: NeRF renders; see §G for a comprehensive TSED analysis.
Scene Completion with Wide Baselines. The quantitative results for the scene completion task are presented in Tab. 2, demonstrating the superiority of our method on all metrics. We also show that our method is less time-consuming than NeRFiller. Moreover, although fitting a NeRF on our inpaintings reduces sharpness, they still remain significantly sharper than NeRFiller’s. To evaluate 3D consistency, we report TSED on two sets of images, (i) the NeRF datasets (i.e. direct outputs of the inpainting models), and (ii) the NeRF renders. For NeRFiller, we use the dataset produced in the latest Dataset Update iteration. As illustrated in Tab. 2, as our dataset used to fit the NeRF is significantly more consistent than that of NeRFiller, the final NeRF achieves a higher consistency score. We qualitatively illustrate this in Fig. 4.
@@
Table 3: Quantitative evaluation of the few-view task. We denote camera parameters as $\Pi$, depth maps as $D$, sharpness (×$\times 10^{-5}$) as $\sigma$, the percentage of consistent image pairs (TSED) at $T_{\text{error}}=2.0$px as $T_{2\text{px}}$, and average run time per scene as $\tau$. see §G for a comprehensive TSED analysis.

### v2

@@ 3.2.2 Uncertainty-aware Multireference Conditioning
Using our cue-based notation, our method inpaints an image $\widehat{I}=f(I,M|A_{\mathcal{R}},\mathcal{G}_{\mathcal{R}})$, based on a reference set $\mathcal{R}$. However, fusing information across references is challenging, especially with conflicting content. We address this by running parallel per-reference diffusion processes and fusing noise estimates at each diffusion step. Ideally, this fusion is geometry-aware (e.g., front-faces of $S_{i}$ provide high-certainty photometric content, while back-faces only upper bound target view depth). Since geometric maps may be slightly misaligned due to errors in estimated geometry, we alter the diffusion model to predict confidence maps for each reference, emulating the aligned geometric masks. In the following, we describe how our inpainting is implemented, including confidence estimation and fusion.
@@
Hierarchical Confidence Estimation. The geometric cues, $\mathcal{G}_{\mathcal{R}}$, indicate the reliable parts of each reference. We utilize three geometric signals. First, the front-facing confidence mask, $\mathbf{C}_{f}$, indicates a pixel is either outside the inpainting mask or guided by a front-facing rendered pixel. In other words, the model has copied the photometric content from the target image itself or the rendered photometric content ($T_{i}(I_{i})$). Second, the back-facing confidence mask, $\mathbf{C}_{b}$, indicates a pixel is geometrically restricted by a rendered back-face, suggesting new geometry to be generated in front of it. Finally, the shadow confidence mask, $\mathbf{C}_{s}$, signals the model’s certainty in trusting photometric information despite potential occlusion (shadow volume). Notice the confidence masks are closely related to the geometric cues $F_{i}$, $B_{i}$, and $C_{i}$. Importantly, though, these cues are often slightly misaligned with the actual target image due to geometry estimation errors. Thus, we instead modify our diffusion model to estimate these confidence masks, i.e., at every diffusion step, our inpainter not only generates a denoising estimate, but also provides an estimate of these three confidence masks. Please see §B.6 for details on supervising the confidence masks, and Fig. 7 for a visualization.
@@
Parallel Diffusion Processing. We now formalize the inpainting process. Given $\mathcal{R}$ and our diffusion model, $f$, we split the inpainter into $n=|\mathcal{R}|$ independent streams. Let $(\varepsilon_{i,t},\mathbf{C}_{f,i,t},\mathbf{C}_{b,i,t},\mathbf{C}_{s,i,t})=f(I,M|A_{i},\mathcal{G}_{i})$, where $\varepsilon_{i,t}$ is the estimated noise for reference $i$, $A_{i}=\{H,T_{i}(I_{i})\}$, and $\mathcal{G}_{i}=\{F_{i},B_{i},\widehat{D}_{i},C_{i}\}$, be the output of the $i$th process at time $t$. Denote $\mathcal{C}_{t}=\{\mathbf{C}_{f,i,t},\mathbf{C}_{b,i,t},\mathbf{C}_{s,i,t}\}_{i}$ as the combined confidence maps. We then fuse the noise estimates, $\mathcal{E}_{t}=\{\varepsilon_{1},\ldots,\varepsilon_{n}\}$, to obtain a fused estimate, $\varepsilon_{t}=\Gamma(\mathcal{E}_{t},\mathcal{C}_{t})$, using our fusion operator, which follows a simple rule-based hierarchy, described in §3.3.1. Before the next timestep, $t-1$, $\varepsilon_{t}$ is used to update the noisy latent $z_{t}$, which is shared across reference streams. This fuses multiview information in the learned, generative space of the diffusion noise estimate, rather than in pixel space.
@@ 3.2.3 Multiview-aware Training
We initialize our geometry-aware inpainter with Stable Diffusion v2, fine-tuned for inpainting [55, 63]. Following prior work [4], we condition on cues by adding zero-initialized channels to the first convolutional layer of the UNet [56]. We also modify the UNet to output confidence masks alongside noise estimates at each timestep. For training, one can utilize a multiview dataset (real or synthetic); however, to ensure greater data diversity, we also synthesize data from single-view images, via monocular depth estimation, perturbing the virtual camera, and rendering the starting image as a reference. To simulate geometry estimation errors, we artificially perturb the reference mesh; in that case, supervision for the confidence masks can be generated via the unperturbed mesh; see §B.6 for additional details.
@@ 3.3 Autoregressive Scene Inpainting
With our geometry estimator (§3.1) and geometry-aware inpainter (§3.2), we can now iteratively inpaint the entire scene. To begin, we initialize the scene geometry by computing multi-view metric depth maps of the incomplete input views. We also initialize the “autoregressive set” with the incomplete input images, which will be autoregressively inpainted. A random view is then selected to start the inpainting. Each iteration consists of three steps: (i) inpainting a subset of not-yet-inpainted images, (ii) updating the autoregressive set and the scene geometry with the inpainted images, and (iii) selecting a subset of images to be inpainted at the next autoregressive iteration. A high-level illustration of these steps and a step-by-step example are provided in Fig. 2 (b) and §F.3, respectively.
@@ 3.3.1 Geometry-aware Inpainting
For each target view, we select a subset of reference views from the autoregressive set, prioritizing those already inpainted. Following §3.2, we render appearance and geometric cues from all reference views, run parallel diffusion processes, and fuse noise estimates at each step via the predicted confidences. The fusion, $\Gamma$, follows a four-level confidence hierarchy: (i) front-face confidence, (ii) back-face confidence, (iii) shadow confidence, and (iv) no confidence. For each patch at each level, we select the noise estimate from the closest camera among the views at the same level. This ensures fusion of the most reliable reference information during denoising; see §B.5 for additional details.
@@ 3.3.2 Autoregressive Set and Scene Geometry Update
In this step, we replace the target views in the autoregressive set with their inpainted versions and update their geometry using the inpainted images, as detailed in §B.1.
@@ 3.3.3 Selecting the Next Images to Inpaint
We employ a two-stage strategy for the autoregressive process. First, we inpaint a wide-baseline subset of the scene, one by one, to generate the missing content of the scene. Then, we propagate the generated content to the remaining views simultaneously. At the start of the first stage, we use a greedy min-max approach to select the wide-baseline subset. Beginning with the first view, we iteratively add the view that maximizes the minimum distance to the existing subset. Once selected, we order them to minimize distance to prior views (see §B.7 for details). The wide-baseline stage follows the sorted order, with each target view conditioned on the entire autoregressive set, inpainted or not. In the propagation step, remaining views are conditioned only on the inpainted wide-baseline images.
@@ 4.1 Implementation Details
We use PyTorch3D [52] for mesh rendering. For training, we use a mixture of two datasets, MS COCO [32] and Google Scanned Objects [15]. For MS COCO, we synthesize the reference-based geometric cues using DepthAnything V2 [79]. Training and evaluation is performed on NVIDIA L40 GPUs (16 for training, one for inference). Additional implementation details are provided in §C.
@@ 4.2 Evaluation Protocol
Datasets. To evaluate our method for narrow-baseline inpainting, we use the SPIn-NeRF [43] dataset, a widely used benchmark for object removal in front-facing real-world scenes. This dataset contains 10 scenes, each with 60 images featuring the object to be removed and corresponding masks, along with 40 images without the object as ground truth for evaluation. In addition, we also use the scene-centric portion of the NeRFiller [75] dataset to further investigate the performance of our method on more complex scenes, particularly those with larger baselines. This dataset serves as a benchmark for the scene completion task. For the few-view inpainting task, we use SPIn-NeRF and the scene-centric portion of NeRFiller, resulting in a total of 15 scenes. For each scene, we uniformly sample eight subsets of two views and eight subsets of three views. This yields a total of 240 few-view sets. Please see §D for details.
@@
Baselines. For object removal, we compare our method to state-of-the-art approaches on the SPIn-NeRF dataset, specifically: SPIn-NeRF [43], which inpaints images and depth maps to supervise NeRF fitting; Inpaint3D [49], which uses SDS [48] to inpaint a NeRF with a diffusion prior; RefFusion [45], which adapts an inpainting diffusion model to a reference image and uses SDS to inpaint a 3DGS [26] with the reference-adapted diffusion model; InFusion [35], which inpaints a 3DGS by inpainting the depth map of an inpainted reference image; MVIP-NeRF [8], which uses SDS to inpaint a NeRF by inpainting the rendered images and normal maps via a diffusion prior, and MALD-NeRF [31], which inpaints a NeRF by performing masked adversarial training for per-scene customization of a diffusion model. InFusion [35] also evaluates Gaussian Grouping [80] on SPIn-NeRF, a 3DGS-based segmentation method that does not directly evaluate on SPIn-NeRF itself. For scene completion, we compare our method to Stable Diffusion [55], which is the naive baseline of independent 2D inpainting, and NeRFiller [75], which alternates between editing input images, and updating the 3D representation encompassed by a NeRF to fuse the edited images, this process is collectively called iterative dataset update. Finally, for the few-view task, we compare our method to SPIn-NeRF [43] and NeRFiller [75]. For this task, we use NeRFiller directly. However, since SPIn-NeRF relies on supervision via COLMAP’s [58] sparse depth, we disable this supervision for the few-view task, as scenes from the NeRFiller dataset lack COLMAP information.
@@
Figure 3: Qualitative object removal comparisons on the SPIn-NeRF dataset. Notice other methods produce blurry regions (e.g., bench end) due to multiview inconsistencies, while ours preserves sharpness and visual plausibility. Please zoom in for details.
@@
Metrics. For the SPIn-NeRF dataset, we use the same evaluation protocol as reported in the original paper [43]. There is no publicly available evaluation code provided with the SPIn-NeRF dataset; we confirmed the details of the protocol with the authors. Specifically, we compute LPIPS [86] (with VGG-16 [61]) and FID [19] between the inpainted images and the ground-truth images, cropped by the inpainting mask’s bounding box. The bounding box’s size is increased by 10% before cropping, uniformly in each direction. We additionally assess the sharpness of the inpainted images within the inpainting mask using the Laplacian variance [47]. As our inpainting method does not explicitly enforce 3D consistency via a 3D radiance field, we evaluate our consistency using the TSED metric [81]. Here, our feature correspondences are limited to those inside the inpainting masks. On the SPIn-NeRF dataset, as the scenes have very small baselines, correspondences are considered across all possible view pairings.
@@
Finally, for the few-view inpainting task, as there is no ground truth available, we only compute sharpness and MUSIQ to evaluate image quality. We also compute Corrs and TSED on all possible image pairs in the few-view set. All metrics are computed only inside the bounding box around the inpainting mask for the few-view inpainting task. §E provides additional details on computing TSED.
@@ 4.3 Results
Object Removal on Narrow-baseline Scenes. Due to the narrow baseline, we perform single-reference inpainting for this task. As our method does not rely on fitting a NeRF on the training views, and the evaluations are performed on the test views, we follow SPIn-NeRF’s [43] procedure to evaluate image inpainters, e.g., LaMa [64]. Specifically, we first fit a NeRF on the training views and render the test views, which will now contain the unwanted object. The rendered test views are then used as inputs to our inpainting method. To evaluate SPIn-NeRF, InFusion [^1] The reported results for InFusion (arXiv-only preprint) are not qualitatively or quantitatively reproducible (others report similar issues on GitHub)., and MVIP-NeRF, we run their official code to reproduce the results. Since RefFusion and Inpaint3D do not provide publicly available code, we report the numbers provided by the papers. For MALD-NeRF, we evaluate their publicly available inpainted images for the SPIn-NeRF dataset. As shown in Tab. 1, our method outperforms all baselines on the SPIn-NeRF benchmark. We obtain comparable sharpness and LPIPS to MALD-NeRF, while significantly outperforming it on other metrics (sharpness and TSED). Further, we demonstrate the ability to handle sparse-view inpainting (see below), which cannot be easily handled by NeRF-based approaches. Please see §F.4 for further analysis. Our method is also efficient, achieving faster scene inpainting compared to other methods. Furthermore, the TSED results show that despite other methods utilizing an explicit or implicit radiance field to enforce 3D consistency, our method achieves a higher consistency score, primarily due to blurry outputs. Please see §G for a comprehensive TSED analysis. Finally, we present a set of qualitative results in Fig. 3, showing our method produces sharper images than baselines, indicating the efficacy of our cross-view fusion, which operates in a learned space, rather than pixel space.
@@
Table 1: Quantitative evaluation of the object removal task on SPIn-NeRF dataset. We denote sharpness (×$\times 10^{-5}$) as $\sigma$, the percentage of consistent image pairs (TSED) at $T_{\text{error}}=2.0$px as $T_{2\text{px}}$, and average run time per scene as $\tau$. The first two sections report numbers from [49] and [35], respectively, possibly using different evaluation code than SPIn-NeRF. The last section ensures direct comparability with consistent evaluation code; see §G for a comprehensive TSED analysis.
@@
Table 2: Evaluation of scene completion on the NeRFiller dataset. We denote sharpness (×$\times 10^{-5}$) as $\sigma$, the percentage of consistent image pairs (TSED) at $T_{\text{error}}=2.0$px as $T_{2\text{px}}$, average run time per scene as $\tau$, and independent 2D inpainting as “2D”. $\cdot^{D}$: direct outputs of the inpainting model - $\cdot^{N}$: NeRF renders; see §G for a comprehensive TSED analysis.
Figure 4: Qualitative scene completion comparisons on the NeRFiller dataset. NeRFiller can converge to blurry content, due to mixing divergent views, while ours generates and then propagates sharp content (e.g., see details in the backpack or window glass in the zoomed patches). Each view in a scene is denoted by $v_{i}$, where $i$ is the view index.
Scene Completion with Wide Baselines. The quantitative results for the scene completion task are presented in Tab. 2, demonstrating the superiority of our method on all metrics. We also show that our method is less time-consuming than NeRFiller. Moreover, although fitting a NeRF on our inpaintings reduces sharpness, they still remain significantly sharper than NeRFiller’s. We qualitatively illustrate this in Fig. 4. To evaluate 3D consistency, we report TSED on two sets of images, (i) the NeRF datasets (i.e., direct outputs of the inpainting models), and (ii) the NeRF renders. For NeRFiller, we use the dataset produced in the latest Dataset Update iteration. As illustrated in Tab. 2, as our dataset used to fit the NeRF is significantly more consistent than that of NeRFiller, the final NeRF achieves a higher consistency score. We also significantly outperform the naive 2D-only baseline (independent inpaintings; see §H.1 for details).
@@
Table 3: Quantitative evaluation of the few-view task. We denote camera parameters as $\Pi$, depth maps as $D$, sharpness (×$\times 10^{-5}$) as $\sigma$, the percentage of consistent image pairs (TSED) at $T_{\text{error}}=2.0$px as $T_{2\text{px}}$, and average run time per scene as $\tau$; see §G for a comprehensive TSED analysis.